# Regularization by regular noise: a numerical result

### v1

@@ 3.1. Stochastic Sewing Lemma
[Ger23, Lemma 2.2] Let $0\leqslant S<T\leqslant 1,p\in[2,\infty),M\geqslant 0$ and let $\left(A_{s,t}\right)_{(s,t)\in[S,T]_{M}^{2}}$ be a family of random variables in $L^{p}\left(\Omega,\mathbb{R}^{d}\right)$ such that $A_{s,t}$ is $\mathcal{F}_{t}$-measurable. Suppose that for some $\varepsilon_{1},\varepsilon_{2}>0$ and $C_{1},C_{2}$ the bounds
@@ 3.2. Fractional Brownian motions
Here we recall the following properties concerning $B^{H}$ that have been used heavily in later analysis.
@@
${\mathbb{E}}|B_{t}^{H}-B_{s}^{H}|\leqslant N|t-s|^{H\wedge 1}.$ (3.6)
@@
For $\alpha,\beta\in[0,1],f\in{\mathcal{C}}^{\alpha},t\in(0,1],$ one has the bounds, with some constant $N$ depending only on $H,\alpha,\beta,d$, accordingly
@@
$\leqslant N\|f\|_{{\mathcal{C}}^{\alpha}}\big(t^{H(\alpha-2)}|x_{1}-x_{2}||x_{1}-x_{3}|+t^{H(\alpha-1)}|x_{1}-x_{2}-x_{3}+x_{4}|\big),\forall x_{i}\in{\mathbb{R}}^{d},i=1,\ldots,4;$ (3.8)
$\|\mathcal{P}_{t}^{H}f\|_{{\mathcal{C}}^{\beta}}\leqslant Nt^{H(\alpha-\beta)\wedge 0}\|f\|_{{\mathcal{C}}^{\alpha}};$ (3.9)
$\|(\mathcal{P}_{t}^{H}-\mathcal{P}_{s}^{H})f\|_{{\mathcal{C}}^{\beta}}\leqslant Ns^{H(\alpha-\beta)-2H\delta}|t^{2H}-s^{2H}|^{\delta}\|f\|_{{\mathcal{C}}^{\alpha}},\forall 0\leqslant s\leqslant t\leqslant 1,0<\delta\in\left[\frac{\alpha-\beta}{2},1\right].$ (3.10)
@@
(3.3) are directly from [Ger23, Page 2 (2.8)]. By using properties of Gaussian convolutions, heat kernel bounds and a relation of the form ${\mathcal{P}}_{t}^{H}f(x)=(p_{c(H)t^{2H}}\ast f)(x)$, we get (3.9) from [BDG21, Proposition 3.7 (i)]. For (3.10), it holds from [BDG21, Proposition 3.7 (ii)]. ∎
@@ 4. Strong convergence rate
Assume Assumption˜2.1 holds. Then for all $t>s$ and $p\geqslant 1$ we have
$\|\varphi_{t}-{\mathbb{E}}^{s}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant N|t-s|^{1+\alpha H}\quad a.s.;$ (4.1)
$\|\varphi^{n}_{t}-{\mathbb{E}}^{s}\varphi_{t}^{n}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant N|t-s|^{1+\alpha H}\quad a.s.$ (4.2)
with some constant $N=N(p,d,\alpha,H,\|b\|_{{\mathcal{C}}^{\alpha}})$.
@@
Suppose (4.1) holds for some $m\geqslant 0$ in place of $1+\alpha H.$ This is certainly true for $m=0$ thanks to the fact that $b$ is bounded; we proceed now by induction on $m.$
We apply (2.2) with $X=\varphi_{t},Y=\varphi_{s}+\int_{s}^{t}b({\mathbb{E}}^{s}B^{H}_{r}+{\mathbb{E}}^{s}\varphi_{r})\mathrm{d}r$ to get
$\|\varphi_{t}-{\mathbb{E}}^{s}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant 2\left\|\varphi_{t}-\varphi_{s}-\int_{s}^{t}b({\mathbb{E}}^{s}B_{r}^{H}+{\mathbb{E}}^{s}\varphi_{r})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$=2\left\|\int_{s}^{t}(b(B^{H}_{r}+\varphi_{r})-b({\mathbb{E}}^{s}B_{r}^{H}+{\mathbb{E}}^{s}\varphi_{r}))\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\left\|\int_{s}^{t}(|B^{H}_{r}-{\mathbb{E}}^{s}B^{H}_{r}|^{\alpha}+|\varphi_{r}-{\mathbb{E}}^{s}\varphi_{r}|^{\alpha})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(\big\||B^{H}_{r}-{\mathbb{E}}^{s}B^{H}_{r}|^{\alpha}\big\|_{L^{p}_{\omega}|\mathcal{F}_{s}}+\big\||\varphi_{r}-{\mathbb{E}}^{s}\varphi_{r}|^{\alpha}\big\|_{L^{p}_{\omega}|\mathcal{F}_{s}})\mathrm{d}r.$
Using (3.7) and the induction hypothesis, we get
$\|\varphi_{t}-{\mathbb{E}}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}\leqslant N|t-s|^{(H\alpha)\wedge(m\alpha)+1}\quad a.s.$
We note that $m_{0}=0,m_{i+1}=1+(H\alpha)\wedge(m_{i}\alpha)$ reaches $1+H\alpha$ in finitely many steps, therefore we get (4.1).
Inequality (4.2) is obtained by a similar argument. For fixed $s,$ define $s^{\prime}$ to be the smallest grid point which is bigger or equal to $s,$ that is, $s^{\prime}:=\lceil ns\rceil n^{-1}$. It is crucial to note that $\varphi_{s^{\prime}}^{n}$ is $\mathcal{F}_{s}$-measurable. We proceed by induction as before and assume that (4.2) holds for some $m\geqslant 0.$ If $s\leqslant t<s^{\prime},$ then $\varphi_{t}^{n}$ is $\mathcal{F}_{s}$-measurable. Hence $\varphi^{n}_{t}={\mathbb{E}}^{s}\varphi_{t}^{n}$ and the left-hand side of (4.2) is zero. Therefore it remains to consider the case $t\geqslant s^{\prime}.$ In this case, using again (2.2) with $X=\varphi_{t}^{n},Y=\varphi^{n}_{s^{\prime}}+\int_{s^{\prime}}^{t}b({\mathbb{E}}^{s}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r,$ we deduce
@@
$\leqslant N\left\|\int_{s^{\prime}}^{t}(|B^{H}_{k_{n}(r)}-{\mathbb{E}}^{s}B^{H}_{k_{n}(r)}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s}\varphi^{n}_{k_{n}(r)}|^{\alpha})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}.$
From now on we obtain (4.2) as before. ∎
@@
Let us start with the estimate for $\mathcal{E}^{b,n,1}$.
@@
Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$ we have
$\left\|\mathcal{E}^{b,n,1}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant N\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}|t-s|^{\frac{1}{2}+\varepsilon},\quad\forall(s,t)\in[0,1]_{0}^{2}$ (4.4)
with sufficiently small $\varepsilon>0$ and some constant $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
@@
$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})-b(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})\mathrm{d}r.$
@@
$A_{s,t}$ $=\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})$
$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{r})\mathrm{d}r.$ (4.5)
@@
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-[s-(t-s)])^{-(1-\alpha)H}\|{\mathbb{E}}^{s-(t-s)}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\mathrm{d}r$ (4.6)
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}|t-s|^{1-(1-\alpha)H}.$
Then (3.1) holds with $C_{1}=N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}$ by the fact that $1-(1-\alpha)H>\frac{1}{2}$.
@@
$=$ ${\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{3}}\int_{s_{4}}^{s_{5}}b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-b(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+b(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{2}}\int_{s_{5}}^{s_{6}}b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-b(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+b(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
$=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
@@
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}|r-s_{3}|^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}$ (4.8)
@@
${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|={\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}(\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})|\leqslant{\mathbb{E}}^{s_{1}}|\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\leqslant N|r-s_{1}|^{1+\alpha H}.$ (4.9)
@@
$\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}\leqslant N|r-s_{1}|^{\frac{1}{2}}[\varphi-\varphi^{n}]_{C^{\frac{1}{2}}_{p}},$ (4.11)
@@
$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\leqslant\|\varphi-\varphi^{n}\|_{C_{p}^{0}},$ (4.12)
@@
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}[\varphi-\varphi^{n}]_{C_{p}^{\frac{1}{2}}}\int_{s_{4}}^{s_{5}}(r-s_{1})^{\frac{1}{2}}(r-s_{3})^{-(1-\alpha)H}\mathrm{d}r$ (4.13)
$\quad+\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}\int_{s_{4}}^{s_{5}}(r-s_{1})^{1+\alpha H}(r-s_{3})^{-(2-\alpha)H}\mathrm{d}r$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{\frac{3}{2}-(1-\alpha)H}[\varphi-\varphi^{n}]_{C_{p}^{\frac{1}{2}}}$
$\quad+\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{2+\alpha H-(2-\alpha)H}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}(t-s)^{(\frac{3}{2}-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)}.$
@@
Noticing Assumption˜2.1 implies $(\frac{3}{2}-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)>1$, we conclude (3.2) holds with $C_{2}=N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}$.
@@
${\mathcal{A}}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi_{r})-b(B^{H}_{r}+\varphi_{r}^{n})\mathrm{d}r.$ (4.14)
@@
${\mathcal{A}}_{t}-{\mathcal{A}}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B^{H}_{r}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r$ (4.15)
$\quad-\int_{s}^{t}b(B_{r}^{H}+\varphi_{r}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})\mathrm{d}r$

### v2

@@ 3.1. Stochastic Sewing Lemma
[13, Lemma 2.2] Let $0\leqslant S<T\leqslant 1,p\in[2,\infty),M\geqslant 0$ and let $\left(A_{s,t}\right)_{(s,t)\in[S,T]_{M}^{2}}$ be a family of random variables in $L^{p}\left(\Omega,\mathbb{R}^{d}\right)$ such that $A_{s,t}$ is $\mathcal{F}_{t}$-measurable. Suppose that for some $\varepsilon_{1},\varepsilon_{2}>0$ and $C_{1},C_{2}$ the bounds
@@ 3.2. Fractional Brownian motions
We recall the following properties concerning $B^{H}$ that have been used heavily in later analysis.
@@
${\mathbb{E}}|B_{t}^{H}-B_{s}^{H}|\leqslant C|t-s|^{H\wedge 1}.$ (3.6)
@@
For $\alpha,\beta\in[0,1],f\in{\mathcal{C}}^{\alpha},t\in(0,1],$ one has the bounds, with some constant $C$ depending only on $H,\alpha,\beta,d$, accordingly
@@
$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\big(t^{H(\alpha-2)}|x_{1}-x_{2}||x_{1}-x_{3}|+t^{H(\alpha-1)}|x_{1}-x_{2}-x_{3}+x_{4}|\big),\forall x_{i}\in{\mathbb{R}}^{d},i=1,\ldots,4;$ (3.8)
$\|\mathcal{P}_{t}^{H}f\|_{{\mathcal{C}}^{\beta}}\leqslant Ct^{H(\alpha-\beta)\wedge 0}\|f\|_{{\mathcal{C}}^{\alpha}};$ (3.9)
$\|(\mathcal{P}_{t}^{H}-\mathcal{P}_{s}^{H})f\|_{{\mathcal{C}}^{\beta}}\leqslant Cs^{H(\alpha-\beta)-2H\delta}|t^{2H}-s^{2H}|^{\delta}\|f\|_{{\mathcal{C}}^{\alpha}},\forall 0\leqslant s\leqslant t\leqslant 1,0<\delta\in\left[\frac{\alpha-\beta}{2},1\right].$ (3.10)
@@
(3.3) are directly from [13, Page 2 (2.8)]. By using properties of Gaussian convolutions, heat kernel bounds and a relation of the form ${\mathcal{P}}_{t}^{H}f(x)=(p_{c(H)t^{2H}}\ast f)(x)$, we get (3.9) from [4, Proposition 3.7 (i)]. For (3.10), it holds from [4, Proposition 3.7 (ii)]. ∎
@@ 4. Strong convergence rate
Assume Assumption˜2.1 holds. Then for all $t>s$ and $p\geqslant 1$ we have a.s.
$\|\varphi_{t}-{\mathbb{E}}^{s}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant C\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1+\alpha H};$ (4.1)
$\|\varphi^{n}_{t}-{\mathbb{E}}^{s}\varphi_{t}^{n}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant C\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1+\alpha H}$ (4.2)
with some constant $C=C(p,d,\alpha,H)$.
@@
We only give the proof of (4.2).
For fixed $s,$ define $s^{\prime}$ to be the smallest grid point which is bigger or equal to $s,$ that is, $s^{\prime}:=\lceil ns\rceil n^{-1}$. It is crucial to note that $\varphi_{s^{\prime}}^{n}$ is $\mathcal{F}_{s}$-measurable. Suppose (4.2) holds for some $m\geqslant 0$ in place of $1+\alpha H.$ This is certainly true for $m=0$ thanks to the fact that $b$ is bounded; we proceed now by induction on $m.$ If $s\leqslant t<s^{\prime},$ then $\varphi_{t}^{n}$ is $\mathcal{F}_{s}$-measurable. Hence $\varphi^{n}_{t}={\mathbb{E}}^{s}\varphi_{t}^{n}$ and the left-hand side of (4.2) is zero. Therefore it remains to consider the case $t\geqslant s^{\prime}.$ In this case, using (2.3) with $X=\varphi_{t}^{n},Y=\varphi^{n}_{s^{\prime}}+\int_{s^{\prime}}^{t}b({\mathbb{E}}^{s}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r,$ we deduce
@@
$\leqslant C\|b\|_{{\mathcal{C}}^{\alpha}}\left\|\int_{s^{\prime}}^{t}(|B^{H}_{k_{n}(r)}-{\mathbb{E}}^{s}B^{H}_{k_{n}(r)}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s}\varphi^{n}_{k_{n}(r)}|^{\alpha})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}.$
Using (3.7) and the induction hypothesis, we get
$\|\varphi_{t}-{\mathbb{E}}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}\leqslant C\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{(H\alpha)\wedge(m\alpha)+1}\quad a.s.$
We note that $m_{0}=0,m_{i+1}=1+(H\alpha)\wedge(m_{i}\alpha)$ reaches $1+H\alpha$ in finitely many steps, therefore we get (4.1). ∎
@@
Let us start with the estimate for $\mathcal{E}^{f,n,1}$.
@@
Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$ and $\gamma\in(\frac{1}{2},1+(\alpha-1)H)$ we have
$\left\|\mathcal{E}^{f,n,1}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\gamma}[s,t]}|t-s|^{\gamma+\varepsilon},\quad(s,t)\in[0,1]_{0}^{2},f\in{\mathcal{C}}^{\alpha}$ (4.4)
with sufficiently small $\varepsilon>0$ and some constant $C=C(p,d,\alpha,\gamma,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
@@
$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}f(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})-f(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})\mathrm{d}r.$
@@
$A_{s,t}$ $=\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}f({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})$
$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}f({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{r})\mathrm{d}r.$ (4.5)
@@
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-[s-(t-s)])^{-(1-\alpha)H}\|{\mathbb{E}}^{s-(t-s)}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\mathrm{d}r$ (4.6)
$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}[s,t]}|t-s|^{1-(1-\alpha)H}.$
Then (3.1) holds with $C_{1}=C\|f\|_{{\mathcal{C}}^{\alpha}[s,t]}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}$ by the fact that $1-(1-\alpha)H>\frac{1}{2}$.
@@
$=$ ${\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{3}}\int_{s_{4}}^{s_{5}}f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-f(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+f(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{2}}\int_{s_{5}}^{s_{6}}f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-f(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-f(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+f(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
$=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{2}}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{2}}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
@@
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}|r-s_{3}|^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}$ (4.8)
@@
${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|={\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}(\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})|\leqslant{\mathbb{E}}^{s_{1}}|\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\leqslant C|r-s_{1}|^{1+\alpha H}.$ (4.9)
@@
$\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}\leqslant C|r-s_{1}|^{\gamma}[\varphi-\varphi^{n}]_{C^{\gamma}_{p}[s,t]},$ (4.11)
@@
$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\leqslant\|\varphi-\varphi^{n}\|_{C_{p}^{0}[s,t]},$ (4.12)
@@
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}[\varphi-\varphi^{n}]_{C_{p}^{\gamma}[s,t]}\int_{s_{4}}^{s_{5}}(r-s_{1})^{\gamma}(r-s_{3})^{-(1-\alpha)H}\mathrm{d}r$ (4.13)
$\quad+\|f\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}[s,t]}\int_{s_{4}}^{s_{5}}(r-s_{1})^{1+\alpha H}(r-s_{3})^{-(2-\alpha)H}\mathrm{d}r$
$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}(t-s)^{1+\gamma-(1-\alpha)H}[\varphi-\varphi^{n}]_{C_{p}^{\gamma}[s,t]}$
$\quad+\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{2+\alpha H-(2-\alpha)H}\|\varphi-\varphi^{n}\|_{C_{p}^{0}[s,t]}$
$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\gamma}[s,t]}(t-s)^{(1+\gamma-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)}.$
@@
Noticing Assumption˜2.1 implies $(1+\gamma-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)>1$, we conclude (3.2) holds with $C_{2}=C\|f\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\gamma}[s,t]}$.
@@
${\mathcal{A}}_{t}=\int_{0}^{t}f(B^{H}_{r}+\varphi_{r})-f(B^{H}_{r}+\varphi_{r}^{n})\mathrm{d}r.$ (4.14)
@@
${\mathcal{A}}_{t}-{\mathcal{A}}_{s}-A_{s,t}$ $=\int_{s}^{t}f(B^{H}_{r}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r$ (4.15)
$\quad-\int_{s}^{t}f(B_{r}^{H}+\varphi_{r}^{n})-\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})\mathrm{d}r$